OpenStudy (anonymous):

Help me set this up... The product of two consecutive positive integers is 5 more than their sum. Find the integers: Let x= 1st integer Let x+5 = 2nd integer What next? Stuck

OpenStudy (anonymous):

Since it says product of two CONSECUTIVE positive integers, Let x = 1st integer x + 1 = 2nd integer Let y be the sum of the two integers So the product of the integers is 5 more than the sum So (x)(x+1) = y +5, Can you take it from here?

OpenStudy (anonymous):

ok, let me see

OpenStudy (anonymous):

(x)(x+1)=y+5 x^2+x=y+5 y=x^2+x-5 Ummm, guess not. Stuck again

OpenStudy (anonymous):

You also have another equation, the sum of the two integers

OpenStudy (anonymous):

\[x(x+1)=5+(x+(x+1))\] then \[x^{2}+x=2x+6\] so \[x^{2}-x-6=0\] so \[(x-3)(x+2)=0\]

OpenStudy (anonymous):

but the number must be positive so the answer for x is 3 and x+1=4

OpenStudy (anonymous):

Ok, that makes sense. I understand it this way Teresita. Thanks, xNiker, I got confused with the y

OpenStudy (anonymous):

Your welcome :), sorry for m grammar, I don't speak english

OpenStudy (anonymous):

That's fine.

OpenStudy (anonymous):

Eh, you could've just went y + 5 = x(x+1), y = x^2 +x - 5 y= x + x + 1 Equate the two 2x + 1 = x^2 + x -5 x^2 - x - 6 = 0, then solve